# Advances in the Shannon Capacity of Graphs

### v2

@@ 8 Outlook
We next prove both directions of the equivalence in (A.1).
@@
    $=\Theta(\mathsf{G})^{i+j}\,\Theta(\mathsf{H})^{i+j},$ (A.5)
    thus $\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\leq\Theta(\mathsf{G})^{i}\,\Theta(\mathsf{H})^{j}$, which implies that for all $k\in\mathbb{N}$
@@
By Theorem 2.15, let $(\mathbf{u}_{1},\ldots,\mathbf{u}_{n})$ be an orthonormal representation of $\overline{\mathsf{G}}$ and let $\mathbf{c}$ be a unit vector such that
@@
Likewise, let $(\mathbf{v}_{1},\ldots,\mathbf{v}_{r})$ be an orthonormal representation of $\overline{\mathsf{H}}$ and let $\mathbf{d}$ be a unit vector such that
@@
Assume without loss of generality that the dimensions of $\mathbf{u}_{i},\mathbf{c},\mathbf{v}_{j}$ and $\mathbf{d}$ are identical and equal to $m$ (if the dimensions are distinct, the vectors of the lower dimension can be padded by zeros). Next, let $\mathbf{A}$ be an orthogonal matrix of order $m\times m$ such that ${\mathbf{A}}{\mathbf{d}}={\mathbf{c}}$. Such a matrix $\mathbf{A}$, satisfying ${\mathbf{A}}^{\mathrm{T}}{\mathbf{A}}=\operatorname{\mathbf{I}}_{m}$, exists (e.g., the householder matrix defined as ${\mathbf{A}}=\operatorname{\mathbf{I}}_{m}-\frac{2(\mathbf{c}-\mathbf{d})(\mathbf{c}-\mathbf{d})^{\mathrm{T}}}{\|\mathbf{c}-\mathbf{d}\|^{2}}$ provided that $\mathbf{c}\neq\mathbf{d}$, and ${\mathbf{A}}=\operatorname{\mathbf{I}}_{m}$ if $\mathbf{c}=\mathbf{d}$). Let $w=(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$ be defined as $\mathbf{w}_{i}=\mathbf{A}\mathbf{v}_{i}$ for every $i\in[r]$. Since the pairwise inner products are preserved under an orthogonal (orthonormal) transformation and $(\mathbf{v}_{1},\ldots,\mathbf{v}_{r})$ is an orthonormal representation of $\overline{\mathsf{H}}$, so is $(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$. Likewise, we get
@@
Appendix C The original proof of Theorem 3.4
@@
Let $\mathsf{G}$ be a simple graph, and let $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\mathcal{U}$ be a mapping from the vertices of $\mathsf{G}$ to a subset $\mathcal{A}$ of $\operatorname{\mathsf{V}}(\mathsf{G})$. $f$ is called an adjacency-reducing mapping if for every nonadjacent vertices $u,v\in\operatorname{\mathsf{V}}(\mathsf{G})$, the vertices $f(u)$ and $f(v)$ are nonadjacent as well.
@@
Let $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{k}\}$ be a maximal independent set of vertices in $\mathsf{G}_{1}$. By assumption, $\operatorname{\mathsf{V}}(\mathsf{G}_{1})$ can be partitioned into $k$ cliques. Let $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$ be such cliques. Obviously every $u_{i},u_{j}\in\mathcal{U}$ cannot be in the same clique, because they are nonadjacent. So, every clique in $\mathcal{C}$ contains exactly one vertex from $\mathcal{U}$. For simplicity, assume that $\mathcal{U}$ and $\mathcal{C}$ are ordered in a way that every $j\in[k]$ has $u_{j}\in C_{j}$. Next, define the mapping $f:\operatorname{\mathsf{V}}(\mathsf{G}_{1})\to\mathcal{U}$ as follows. Let $v\in\operatorname{\mathsf{V}}(\mathsf{G}_{1})$ be a vertex, and let $C_{j}$ be the clique that has $v\in C_{j}$, then $f(v)=u_{j}$. If $v_{1}$ and $v_{2}$ are nonadjacent, then they belong to different cliques in $\mathcal{C}$. Thus $f(v_{1})$ and $f(v_{2})$ are mapped to different vertices in $\mathcal{U}$. Since $\mathcal{U}$ is an independent set, $f(v_{1})$ and $f(v_{2})$ are nonadjacent. Thus, $f$ is an adjacency-reducing mapping of $\mathsf{G}_{1}$ into $\mathcal{U}$. Then, by [1, Theorem 4], equality (77) holds. ∎
Appendix D The original proof of Theorem 5.1
@@
Let $\mathsf{G}$ be a universal graph and let $k\in\mathbb{N}$. Then, since $\mathsf{G}$ is universal,
@@
Note that if we assume that the equality $\Theta(\mathsf{K})=\sqrt[k]{\alpha(\mathsf{K}^{k})}$ holds for some $k<k_{0}$, then for every $n\in\mathbb{N}$
@@
Appendix E The original proof of inequality (245)
@@
[7] Let $\mathsf{G}$ be a simple graph with $n$ vertices. Then
@@
In addition, inequality (E.1) holds with equality if $\mathsf{G}$ is self-complementary and vertex-transitive.
@@
Let $\mathcal{A}\cup\mathcal{B}=\{a_{1},\ldots,a_{n},b_{1},\ldots,b_{n}\}$ be the vertex set of $\mathsf{G}+\overline{\mathsf{G}}$, where $\mathcal{A}$ represents the vertices of $\mathsf{G}$ and $\mathcal{B}$ represents the matching vertices of $\overline{\mathsf{G}}$. Thus, if $\{a_{i},a_{j}\}\in\operatorname{\mathsf{E}}(\mathsf{G})$, then $\{b_{i},b_{j}\}\notin\operatorname{\mathsf{E}}(\overline{\mathsf{G}})$. Next, we construct an independent set in the graph $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$ for every $k$, and we use that set to bound the capacity of $\mathsf{G}+\overline{\mathsf{G}}$. For every $k$, define the set $\mathcal{S}$ as the set of vectors $\textbf{v}=(v_{1},v_{2},\ldots,v_{2k})$ that follows the following two rules:
  • $|\{i:v_{i}\in\mathcal{A}\}|=|\{j:v_{j}\in\mathcal{B}\}|$,
  • For every $1\leq i\leq k$, if in v, $a_{r}$ is the $i$-th coordinate from the left that belongs to $\mathcal{A}$, and $b_{s}$ is the $i$-th coordinate from the left that belongs to $\mathcal{B}$, then $r=s$.
Now we prove that $\mathcal{S}$ is an independent set in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. Let u and v be two distinct vectors from $\mathcal{S}$, we consider two cases:
  • Case 1: if there is an index $t$ that has $u_{t}\in\mathcal{A}$ and $v_{t}\in\mathcal{B}$ then obviously the two vertices u and v are not adjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
  • Case 2: if there isn’t such an index, then there are two indices $1\leq i,j\leq k$, and two indices $r,s$ between $1$ and $2k$ that have $u_{r}=a_{i}$, $u_{s}=b_{i}$, and $v_{r}=a_{j}$, $v_{s}=b_{j}$. Since $\{a_{i},a_{j}\}$ is an edge in $\mathsf{G}$ if and only if $\{b_{i},b_{j}\}$ is not an edge in $\overline{\mathsf{G}}$, then the vertices u and v are not adjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. Hence, $\mathcal{S}$ is an independent set in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
The cardinality of $\mathcal{S}$ is $\binom{2k}{k}\,n^{k}$. Indeed, there are $\binom{2k}{k}$ ways to choose the positions of the coordinates from $\mathcal{A}$, which uniquely determines the positions of the coordinates from $\mathcal{B}$. For each such choice, there are $n$ possible values for each of the $k$ coordinates, yielding $|\mathcal{S}|=\binom{2k}{k}\,n^{k}$. Finally, we get
$\alpha((\mathsf{G}+\overline{\mathsf{G}})^{2k})\geq|\mathcal{S}|=\binom{2k}{k}\,n^{k},$ (E.2)
which gives
@@
In addition, from [11, Theorem 12], inequality (E.1) indeed holds with equality when $\mathsf{G}$ is a self-complementary and vertex-transitive graph. ∎

### v3

@@ 8 Outlook
  6. 6.
    This last open issue, while not directly related to the results of this paper, is mentioned here because of its fundamental importance. The random graph $G(n,\tfrac{1}{2})$ is a graph on $n$ vertices in which each pair of distinct vertices is joined by an edge independently with probability $\tfrac{1}{2}$. For sufficiently large $n$, with high probability, the independence number of such a graph is equal to $2\log_{2}n+O(\log\log n)$ [54, Chapter 11], and its Lovász $\vartheta$-function ranges between $\bigl(\tfrac{1}{2}+o(1)\bigr)\sqrt{n}$ and $\bigl(2+o(1)\bigr)\sqrt{n}$ [55]. In contrast, a comparable probabilistic result for the Shannon capacity of $G(n,\tfrac{1}{2})$ is yet unknown. It has been conjectured by Alon that there exists a positive constant $b$, independent of $n$, such that the Shannon capacity of $G(n,\tfrac{1}{2})$ is at most $b\log_{2}n$ almost surely (i.e., with probability tending to 1 as $n\to\infty$) [3, Conjecture 2.4].
This sample of open problems illustrates that, despite the existence of classical results on the Shannon capacity of graphs and a growing body of recent work, many fundamental avenues of research remain open.
@@
Next, we prove both directions of the equivalence in (A.1).
@@
    $=\Theta(\mathsf{G})^{i+j}\,\Theta(\mathsf{H})^{i+j}.$ (A.5)
    Thus, $\Theta(\mathsf{G}^{i}\boxtimes\mathsf{H}^{j})\leq\Theta(\mathsf{G})^{i}\,\Theta(\mathsf{H})^{j}$, which implies that for all $k\in\mathbb{N}$,
@@
By Theorem 2.15, let $(\mathbf{u}_{1},\ldots,\mathbf{u}_{n})$ be an orthonormal representation of $\overline{\mathsf{G}}$, and let $\mathbf{c}$ be a unit vector such that
@@
Likewise, let $(\mathbf{v}_{1},\ldots,\mathbf{v}_{r})$ be an orthonormal representation of $\overline{\mathsf{H}}$, and let $\mathbf{d}$ be a unit vector such that
@@
Assume without loss of generality that the dimensions of $\mathbf{u}_{i},\mathbf{c},\mathbf{v}_{j}$, and $\mathbf{d}$ are identical and equal to $m$ (if the dimensions are distinct, the vectors of the lower dimension can be padded by zeros). Next, let $\mathbf{A}$ be an orthogonal matrix of order $m\times m$ such that ${\mathbf{A}}{\mathbf{d}}={\mathbf{c}}$. Such a matrix $\mathbf{A}$, satisfying ${\mathbf{A}}^{\mathrm{T}}{\mathbf{A}}=\operatorname{\mathbf{I}}_{m}$, exists (e.g., the householder matrix defined as ${\mathbf{A}}=\operatorname{\mathbf{I}}_{m}-\frac{2(\mathbf{c}-\mathbf{d})(\mathbf{c}-\mathbf{d})^{\mathrm{T}}}{\|\mathbf{c}-\mathbf{d}\|^{2}}$ provided that $\mathbf{c}\neq\mathbf{d}$, and ${\mathbf{A}}=\operatorname{\mathbf{I}}_{m}$ if $\mathbf{c}=\mathbf{d}$). Let $w=(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$ be defined as $\mathbf{w}_{i}=\mathbf{A}\mathbf{v}_{i}$ for every $i\in[r]$. Since the pairwise inner products are preserved under an orthogonal (orthonormal) transformation and $(\mathbf{v}_{1},\ldots,\mathbf{v}_{r})$ is an orthonormal representation of $\overline{\mathsf{H}}$, so is $(\mathbf{w}_{1},\ldots,\mathbf{w}_{r})$. Likewise, we get
@@
Appendix C Proof of Theorem 3.4
@@
Let $\mathsf{G}$ be a simple graph, and let $f\colon\operatorname{\mathsf{V}}(\mathsf{G})\to\mathcal{A}$ be a mapping from the vertices of $\mathsf{G}$ to a subset $\mathcal{A}$ of $\operatorname{\mathsf{V}}(\mathsf{G})$. The mapping $f$ is called adjacency-reducing if for every pair of nonadjacent vertices $u,v\in\operatorname{\mathsf{V}}(\mathsf{G})$, the vertices $f(u)$ and $f(v)$ are also nonadjacent.
@@
Let $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{k}\}$ be a maximal independent set of vertices in $\mathsf{G}_{1}$. By assumption, $\operatorname{\mathsf{V}}(\mathsf{G}_{1})$ can be partitioned into $k$ cliques. Let $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$ be such cliques. Obviously, every $u_{i},u_{j}\in\mathcal{U}$ cannot be in the same clique because they are nonadjacent. So, every clique in $\mathcal{C}$ contains exactly one vertex from $\mathcal{U}$. For simplicity, assume that $\mathcal{U}$ and $\mathcal{C}$ are ordered in a way that $u_{j}\in C_{j}$ for every $j\in[k]$. Next, define the mapping $f:\operatorname{\mathsf{V}}(\mathsf{G}_{1})\to\mathcal{U}$ as follows. Let $v\in\operatorname{\mathsf{V}}(\mathsf{G}_{1})$ be a vertex, and let $C_{j}$ be the clique that has $v\in C_{j}$. Then, $f(v)=u_{j}$. If $v_{1}$ and $v_{2}$ are nonadjacent, then they belong to different cliques in $\mathcal{C}$. Thus, $f(v_{1})$ and $f(v_{2})$ are mapped to different vertices in $\mathcal{U}$. Since $\mathcal{U}$ is an independent set, $f(v_{1})$ and $f(v_{2})$ are nonadjacent. Thus, $f$ is an adjacency-reducing mapping of $\mathsf{G}_{1}$ into $\mathcal{U}$. Then, by [1, Theorem 4], equality (77) holds. ∎
Appendix D The original proof from [10] of Theorem 5.1
@@
Let $\mathsf{G}$ be a universal graph, and let $k\in\mathbb{N}$. Then, since $\mathsf{G}$ is universal,
@@
Note that if we assume that the equality $\Theta(\mathsf{K})=\sqrt[k]{\alpha(\mathsf{K}^{k})}$ holds for some $k<k_{0}$, then for every $n\in\mathbb{N}$,
@@
Appendix E The original proof from [7] of inequality (245)
@@
[7] Let $\mathsf{G}$ be a simple graph on $n$ vertices. Then,
@@
Let $\mathcal{A}\cup\mathcal{B}=\{a_{1},\ldots,a_{n},b_{1},\ldots,b_{n}\}$ be the vertex set of $\mathsf{G}+\overline{\mathsf{G}}$, where $\mathcal{A}$ and $\mathcal{B}$ are the vertex sets of a copy of $\mathsf{G}$ and of its complement $\overline{\mathsf{G}}$, respectively. The labeling is chosen so that, for $i\neq j$, $\{a_{i},a_{j}\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ if and only if $\{b_{i},b_{j}\}\notin\operatorname{\mathsf{E}}(\overline{\mathsf{G}})$. Moreover, since $\mathsf{G}+\overline{\mathsf{G}}$ is a disjoint union, there are no edges of the form $\{a_{i},b_{j}\}$. For each $k\in\mathbb{N}$, we construct an independent set in the strong power $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$, which is then used to bound the Shannon capacity of $\mathsf{G}+\overline{\mathsf{G}}$. For every $k$, define $\mathcal{S}_{k}$ as the set of all vectors $\textbf{v}=(v_{1},v_{2},\ldots,v_{2k})$ with coordinates in $\mathcal{A}\cup\mathcal{B}$ that satisfy the following two conditions:
  1. 1.
    $\bigl|\{i:v_{i}\in\mathcal{A}\}\bigr|=\bigl|\{j:v_{j}\in\mathcal{B}\}\bigr|=k$,
  2. 2.
    For every $i\in[k]$, if $a_{r}$ is the $i$th coordinate of v (from left to right) belonging to $\mathcal{A}$, and $b_{s}$ is the $i$th coordinate of v belonging to $\mathcal{B}$, then $r=s$.
Next, we prove that $\mathcal{S}_{k}$ is an independent set in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. Let u and v be two distinct vectors in $\mathcal{S}_{k}$. We consider the following two cases:
  • Case 1: If there exists an index $t$ such that $u_{t}\in\mathcal{A}$ and $v_{t}\in\mathcal{B}$ (or vice versa), then the vertices u and v are clearly nonadjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
  • Case 2: If there is no such an index, then there exist two indices $1\leq i,j\leq k$, and two positions $1\leq r,s\leq 2k$ such that $u_{r}=a_{i}$, $u_{s}=b_{i}$, $v_{r}=a_{j}$, and $v_{s}=b_{j}$. Since $\{a_{i},a_{j}\}\in\operatorname{\mathsf{E}}(\mathsf{G})$ if and only if $\{b_{i},b_{j}\}\notin\operatorname{\mathsf{E}}(\overline{\mathsf{G}})$, the vertices u and v are again nonadjacent in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$.
Hence, $\mathcal{S}_{k}$ is an independent set in $(\mathsf{G}+\overline{\mathsf{G}})^{2k}$. The cardinality of $\mathcal{S}_{k}$ is $\binom{2k}{k}\,n^{k}$. Indeed, there are $\binom{2k}{k}$ ways to choose the positions of the coordinates from $\mathcal{A}$, which uniquely determines the positions of the coordinates from $\mathcal{B}$. For each such choice, there are $n$ possible values for each of the $k$ coordinates, yielding $|\mathcal{S}_{k}|=\binom{2k}{k}\,n^{k}$. Consequently,
$\alpha((\mathsf{G}+\overline{\mathsf{G}})^{2k})\geq|\mathcal{S}_{k}|=\binom{2k}{k}\,n^{k},$ (E.8)
which yields
@@
Remark E.1.
It is noted in [7] that the inequality $\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq\sqrt{2n}$, which is a loosened version of inequality (E.7), easily follows for every simple graph $\mathsf{G}$ on $n$ vertices. This holds since
$\bigl\{(a_{i},b_{i})\bigr\}_{i\in[n]}\bigcup\bigl\{(b_{i},a_{i})\bigr\}_{i\in[n]}$
is an independent set of the square $(\mathsf{G}+\overline{\mathsf{G}})^{2}$, consisting of $2n$ elements. Furthermore, it is also noted in [7] that inequality (E.7) is easily obtained for every self-complementary and vertex-transitive graph $\mathsf{G}$. Indeed, by [11, Theorem 12], $\Theta(\mathsf{G})=\sqrt{n}=\Theta(\overline{\mathsf{G}})$ for every such graph on $n$ vertices, so it follows from Shannon’s inequality (31) that $\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq 2\sqrt{n}$. Our strengthened result in Corollary 7.1 asserts that inequality (E.7) holds with equality for all self-complementary vertex-transitive graphs, as well as for some additional graph families.